OpenStudy (anonymous):

101, 102, 100, 100, 110, 109, 109, 108, 109 d. Find the variance of the data. e. Find the standard deviation of the data.

OpenStudy (anonymous):

ill help you

OpenStudy (anonymous):

thank you @prowrestler

OpenStudy (john_es):

For the variance of the data you need first the mean value of the data, \[\overline{x}=(101+ 102+ 100+ 100+ 110+ 109+ 109+ 108+ 109)/9\] As you know, the sum of all values, divided by the number of values. Once you get this, the variance can be calculated with this formula, \[Var=\frac{\sum_i(x_i-\overline{x})^2}{N}\] Where x_i is each data, N the number of data, and the overlined x is the mean. For the standard deviation, \[\sigma=\sqrt{Var}\]

OpenStudy (anonymous):

I dont understand... @John_ES

OpenStudy (john_es):

First, do you understand how to calculate the mean?

OpenStudy (john_es):

When you calculate the mean value of your data, you will obtain this, \[\overline{x}=(101 + 102 + 100 + 100 + 110 + 109 + 109 + 108 + 109)/9=316/3\\ \overline{x}\approx105.3\]

OpenStudy (anonymous):

yes, but I dont know how to use this Var=∑i(xi−x¯)2N @John_ES

OpenStudy (john_es):

Ok, I will expand the formula using your data

OpenStudy (john_es):

\[Var=((101-105.3)^2+(102-105.3)^2+2\cdot(100-105.3)^2+\\ +(108-105.3)^2+3\cdot(109-105.3)^2+(110-105.3)^2)/9\]

OpenStudy (john_es):

Ok, then, \[Var\approx17.33\] And \[\sigma\approx\sqrt{Var}=4.16\]

OpenStudy (john_es):

Only for checking.